∫ x tan² 3x² dx


1
Expert's answer
2021-04-15T07:52:09-0400

"\displaystyle\int x \tan^2 (3x^2) dx"


Note that if

"d(3x^2)=(3x^2)'dx=6xdx"

then

"xdx=\dfrac{1}{6}\cdot d(3x^2)"


"\displaystyle\int x\cdot \tan^2 (3x^2) dx="


"\displaystyle\int \tan^2 (3x^2)\cdot x dx="

"\displaystyle\int \tan^2 (3x^2)\cdot \dfrac{1}{6}\cdot d(3x^2)="


"\dfrac{1}{6}\cdot\displaystyle\int \tan^2 (3x^2)\cdot d(3x^2)"


Further substitute

"\tan (3x^2)=t"

Hence

"3x^2=\arctan(t)"

"d(3x^2)=d(\arctan(t))=(\arctan(t))'dt=\dfrac{1}{t^2+1}dt"



Continue solving

"\dfrac{1}{6}\cdot\displaystyle\int \tan^2 (3x^2)\cdot d(3x^2)="


"\dfrac{1}{6}\cdot\displaystyle\int t^2 \cdot \dfrac{1}{t^2+1}dt="


"\dfrac{1}{6}\cdot\displaystyle\int \dfrac{t^2}{t^2+1}dt="


"\dfrac{1}{6}\cdot\displaystyle\int \dfrac{t^2+1-1}{t^2+1}dt="


"\dfrac{1}{6}\cdot\displaystyle\int \left(\dfrac{t^2+1}{t^2+1}-\dfrac{1}{t^2+1}\right)dt="


"\dfrac{1}{6}\cdot\displaystyle\int \left(1-\dfrac{1}{t^2+1}\right)dt="


"\dfrac{1}{6}\cdot\displaystyle \left(t-\arctan(t)\right)+c=" <return to variable "x": "t=\tan (3x^2)" >


"\dfrac{1}{6}\cdot\displaystyle \left(\tan (3x^2)-\arctan(\tan (3x^2))\right)+c="


"\dfrac{1}{6}\cdot\displaystyle \left(\tan (3x^2)-3x^2\right)+c="


"\dfrac{1}{6}\tan (3x^2)-\dfrac{1}{6}\cdot3x^2+c="


"\dfrac{1}{6}\tan (3x^2)-\dfrac{x^2}{2} +c"